A large reservoir is drained through a pipe with an internal diameter of 15 cm.
If the opening of the pipe is 17 m below the surface of the water in the reservoir then at what speed is the water exiting the pipe?
(treat water as an incompressible non-viscous fluid)
1
Expert's answer
2020-04-08T10:53:58-0400

As per the given question,

Internal diameter of the pipe "(d) =15cm"

Opening of the pipe is 17m below, let the velocity of the water at the below is v

Now, applying the bernaulli theorem,

"\rho gh+\dfrac{\rho v^2}{2}=constant"

"v=\sqrt{2gh}"

"v=\sqrt{2\times 9.8\times 17}=18.25m/sec"